# Distributional Finite Element curl div Complexes and Application to Quad Curl Problems

### v3

@@ 3.3. Characterization of null spaces
The first Poincaré inequality (20) can be found in [19, Lemma 3.4 and Theorem 3.6] and [35, Lemma 7.20]. Consequently $({\rm curl\,}\cdot,{\rm curl\,}\cdot)$ is an inner product on $K_{h}^{c}$ and the operator $({\rm curl\,}{\rm curl\,})_{h}$ is isomorphic.
@@ 3.4. Helmholtz decompositions
4. Mixed finite element methods of the quad-curl problem
@@ 4.1. Distributional mixed formulation
(33) $(\boldsymbol{\sigma},\boldsymbol{\tau})+b(\boldsymbol{\tau},\psi;\boldsymbol{u})$ $=0,$ $\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega),$
(34) $b(\boldsymbol{\sigma},\phi;\boldsymbol{v})$ $=-\langle\boldsymbol{f},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega),$
@@
Combining (35), (37) and the Babuška-Brezzi theory [5] yileds the wellposedness of the mixed formulation (33)-(34).
@@
The mixed formulation (33)-(34) is well-posed. Namely for any $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$, there exists a unique solution $(\boldsymbol{\sigma},\boldsymbol{u},\phi)$ to (33)-(34). Furthermore we have $\phi=0$, and the stability
@@
Combine the inf-sup condition (35) and the coercivity (37) to get (38) and the well-posedness of the mixed formulation (33)-(34). By choosing $\boldsymbol{v}={\rm grad\,}\phi$ in (34), we get $\phi=0$ from $\operatorname{div}\boldsymbol{f}=0$. ∎
Remark 4.4.
The quad-curl problem (31) is equivalent to the Stokes equation with ${\rm curl\,}\boldsymbol{u}$ being the velocity. The mixed formulation (33)-(34) is equivalent to the distributional formulation of Stokes equation [21]: find $\boldsymbol{\sigma}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T})$, $\boldsymbol{u}\in H_{0}(\operatorname{div},\Omega)$ and $p\in L_{0}^{2}(\Omega)$ such that
$(\boldsymbol{\sigma},\boldsymbol{\tau})+\langle\operatorname{div}\boldsymbol{\tau},\boldsymbol{u}\rangle-(\operatorname{div}\boldsymbol{u},q)$ $=0,$ $\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),q\in L_{0}^{2}(\Omega),$
$\langle\operatorname{div}\boldsymbol{\sigma},\boldsymbol{v}\rangle-(\operatorname{div}\boldsymbol{v},p)$ $=-\langle\boldsymbol{g},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in H_{0}(\operatorname{div},\Omega),$
[…]
$\langle\operatorname{div}\boldsymbol{\tau},\boldsymbol{v}\rangle=\sum_{T\in\mathcal{T}_{h}}(\operatorname{div}\boldsymbol{\tau},\boldsymbol{v})_{T}-\sum_{T\in\mathcal{T}_{h}}(\boldsymbol{n}^{\intercal}\boldsymbol{\tau}\boldsymbol{n},\boldsymbol{v}\cdot\boldsymbol{n})_{\partial T}.$
The tangential-normal continuous finite element space $\Sigma_{k,h}^{\rm tn}$ can be used to discretize this formulation.
4.2. Distributional mixed finite element methods
@@
(39) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},$
(40) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-\langle\boldsymbol{f},\boldsymbol{v}_{h}\rangle,$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}.$
@@
(41) $\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|\psi_{h}|_{1}}.$
@@
Applying the Babuška-Brezzi theory [5], from the discrete inf-sup conditions (41) and (43), and the discrete coercivity (45), we achieve the well-posedness of the mixed finite element method (39)-(40).
@@
 The distributional mixed finite element method (39)-(40) is well-posed. We have the discrete stability results
@@
By choosing $\boldsymbol{v}_{h}={\rm grad\,}\phi_{h}$ in (40), we get $\phi_{h}=0$ from $\operatorname{div}\boldsymbol{f}=0$.
@@ 4.3. Error analysis
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (33)-(34) and the mixed finite element method (39)-(40) respectively. Assume $\boldsymbol{\sigma}\in\Sigma^{\rm tn}$, and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{1}(\Omega;\mathbb{R}^{3})$. Then
@@
Subtract (39)-(40) from (33)-(34) and use (22) and (23) to get error equations
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (33)-(34) and the mixed finite element method (39)-(40) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
@@
where $\widetilde{\boldsymbol{\sigma}}:={\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}}$. When $\Omega$ is convex, regularity (52) has been obtained in [29, Lemma A.1].
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (33)-(34) and the mixed finite element method (39)-(40) respectively. Assume the regularity condition (52) holds, $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{3})$, $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
@@
which together with (33) and (39) gives
@@
On the other side, it follows from (34), (40) and (23) that
@@
We perform numerical experiments to support the theoretical results of the distributional mixed method (39)-(40). Let $\Omega=(0,1)^{3}$. Choose the function $\boldsymbol{f}$ in (31) such that the exact solution of (31) is
$\boldsymbol{u}={\rm curl\,}\begin{pmatrix}0\\ 0\\ \sin^{2}(\pi x)\sin^{2}(\pi y)\sin^{2}(\pi z)\end{pmatrix},$
and let $\boldsymbol{\sigma}:={\rm grad\,}{\rm curl\,}\boldsymbol{u}$. We take uniform triangulations on $\Omega$.
Numerical results of $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|$, $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ of the distributional mixed method (39)-(40) with $k=1$ and $\ell=0,1$ are shown in Table 1. It is observed that $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h)$, and $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\mathcal{O}(h)$ numerically, which coincide with the theoretical error estimates in (49) and (50). The error $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|=\mathcal{O}(h^{\ell+1})$ is also observed but not included in (49) and (50).
Table 1. Errors for the distributional mixed finite element method (39)-(40) with $k=1$.
$h$ $2^{-1}$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ 1.03E+02 9.46E+01 5.44E+01 2.83E+01 1.43E+01
order -- 0.12 0.79 0.94 0.98
$\|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 9.50E+00 8.67E+00 3.74E+00 1.73E+00 8.46E-01
order -- 0.13 1.21 1.10 1.03
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 1.04E+02 1.04E+02 1.04E+02 1.04E+02 1.04E+02
order -- 0 0 0 0
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=0$ 9.86E-01 8.61E-01 3.50E-01 1.60E-01 7.77E-02
order -- 0.19 1.29 1.13 1.03
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=1$ 9.85E-01 5.41E-01 1.47E-01 3.79E-02 9.56E-03
order -- 0.86 1.87 1.95 1.98
@@ 4.4. Post-processing
We will construct a new superconvergent approximation to $\boldsymbol{u}$ in virtue of the optimal result of $\boldsymbol{\sigma}$ in (49) and the superconvergent result (53).
Introduce discrete space
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (33)-(34) and the mixed finite element method (39)-(40) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k+1}(\Omega;\mathbb{R}^{3})$. Then
@@
$h$ $2^{-1}$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
$\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 9.34E+00 5.32E+00 1.58E+00 4.20E-01 1.07E-01
order -- 0.82 1.75 1.91 1.97
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 1.03E+02 9.46E+01 5.44E+01 2.83E+01 1.43E+01
order -- 0.12 0.79 0.94 0.98
Numerical results of the post-processing $\boldsymbol{u}_{h}^{\ast}$ for $k=1$ and $\ell=0,1$ are listed in Table 2. We can see that $\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h^{2})$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h)$ numerically, which are one order higher than $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ respectively. Indeed, $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$.
5. Hybridization
In this section we will hybridize the distributional mixed finite element method (39)-(40) following the framework in [1]. We introduce a Lagrange multiplier for the tangential-normal continuity of $\boldsymbol{\sigma}$ which can be treat as an approximation of $(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})|_{\mathcal{F}_{h}}$. As $\boldsymbol{\sigma}$ is discontinuous, it can be eliminated element-wise and the size of the resulting linear system is reduced, which is easier to solve than the saddle point system obtained by the mixed method (39)-(40). With the hybridized method, we can also establish the equivalence of the mixed method (39)-(40) to a weak Galerkin method without stabilization and nonconforming finite element methods in [29, 47].

### v4

@@ 3.3. Characterization of null spaces
The first Poincaré inequality (21) can be found in [20, Lemma 3.4 and Theorem 3.6] and [36, Lemma 7.20]. Consequently $({\rm curl\,}\cdot,{\rm curl\,}\cdot)$ is an inner product on $K_{h}^{c}$ and the operator $({\rm curl\,}{\rm curl\,})_{h}$ is isomorphic.
@@ 3.4. Helmholtz decompositions
4. Mixed finite element method of the quad-curl problem
@@ 4.1. Distributional mixed formulation
(34a) $(\boldsymbol{\sigma},\boldsymbol{\tau})+b(\boldsymbol{\tau},\psi;\boldsymbol{u})$ $=0,$ $\boldsymbol{\tau}\in H^{-1}(\operatorname{curl}\operatorname{div},\Omega;\mathbb{T}),\psi\in H_{0}^{1}(\Omega),$
(34b) $b(\boldsymbol{\sigma},\phi;\boldsymbol{v})$ $=-\langle\boldsymbol{f},\boldsymbol{v}\rangle,$ $\boldsymbol{v}\in H_{0}({\rm curl\,},\Omega),$
@@
Combining (35), (37) and the Babuška-Brezzi theory [6] yileds the wellposedness of the mixed formulation (34a)-(34b).
@@
The mixed formulation (34a)-(34b) is well-posed. Namely for any $\boldsymbol{f}\in H^{-1}(\operatorname{div},\Omega)\cap\ker(\operatorname{div})$, there exists a unique solution $(\boldsymbol{\sigma},\boldsymbol{u},\phi)$ to (34a)-(34b). Furthermore we have $\phi=0$, and the stability
@@
Combine the inf-sup condition (35) and the coercivity (37) to get (38) and the well-posedness of the mixed formulation (34a)-(34b). By choosing $\boldsymbol{v}={\rm grad\,}\phi$ in (34b), we get $\phi=0$ from $\operatorname{div}\boldsymbol{f}=0$. ∎
4.2. Distributional mixed finite element method
@@
(39a) $(\boldsymbol{\sigma}_{h},\boldsymbol{\tau}_{h})+b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{u}_{h})$ $=0,$ $\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}},$
(39b) $b_{h}(\boldsymbol{\sigma}_{h},\phi_{h};\boldsymbol{v}_{h})$ $=-\langle\boldsymbol{f},\boldsymbol{v}_{h}\rangle,$ $\boldsymbol{v}_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm curl\,}}.$
@@
(40) $\|\boldsymbol{v}_{h}\|_{H({\rm curl\,})}\lesssim\sup_{\boldsymbol{\tau}_{h}\in{\Sigma}^{\rm tn}_{h},\psi_{h}\in\mathring{\mathbb{V}}_{h}^{{\rm grad\,}}}\frac{b_{h}(\boldsymbol{\tau}_{h},\psi_{h};\boldsymbol{v}_{h})}{\|\boldsymbol{\tau}_{h}\|_{H^{-1}(({\rm curl\,}\operatorname{div})_{h})}+|\psi_{h}|_{1}}.$
@@
Applying the Babuška-Brezzi theory [6], from the discrete inf-sup conditions (40) and (42), and the discrete coercivity (44), we achieve the well-posedness of the mixed finite element method (39a)-(39b).
@@
 The distributional mixed finite element method (39a)-(39b) for the quad-curl problem is well-posed. We have the discrete stability results
@@
By choosing $\boldsymbol{v}_{h}={\rm grad\,}\phi_{h}$ in (39b), we get $\phi_{h}=0$ from $\operatorname{div}\boldsymbol{f}=0$.
@@ 4.3. Error analysis
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in\Sigma^{\rm tn}$, and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{1}(\Omega;\mathbb{R}^{3})$. Then
@@
Subtract (39a)-(39b) from (34a)-(34b) and use (23) and (24) to get error equations
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
@@
where $\widetilde{\boldsymbol{\sigma}}:={\rm grad\,}{\rm curl\,}\widetilde{\boldsymbol{u}}$. When $\Omega$ is convex, regularity (51) has been obtained in [30, Lemma A.1].
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume the regularity condition (51) holds, $\boldsymbol{f}\in L^{2}(\Omega;\mathbb{R}^{3})$, $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k}(\Omega;\mathbb{R}^{3})$. Then
@@
which together with (34a) and (39a) gives
@@
On the other side, it follows from (34b), (39b) and (24) that
@@
We perform numerical experiments to support the theoretical results of the distributional mixed method (39a)-(39b). Let $\Omega=(0,1)^{3}$. Choose the function $\boldsymbol{f}$ in (32) such that the exact solution of (32) is
$\boldsymbol{u}={\rm curl\,}\begin{pmatrix}\sin^{3}(\pi x)\sin^{3}(\pi y)\sin^{3}(\pi z)\\ \sin^{3}(\pi x)\sin^{3}(\pi y)\sin^{3}(\pi z)\\ 0\end{pmatrix},$
and let $\boldsymbol{\sigma}:={\rm grad\,}{\rm curl\,}\boldsymbol{u}$. To break the symmetry, the initial unstructured mesh of $\Omega$ is shown in Fig. 4, which is a perturbation of the uniform mesh with $h=1/2$. Then we take uniform refinement of this initial unstructured triangulation.
Figure 4. An initial perturbed mesh of the uniform mesh with $h=1/2$.
We will implement the hybridized version of (39a)-(39b); see (60a)-(60b). Therefore the number of DoFs is $\dim\mathring{\mathbb{V}}^{{\rm curl\,}}_{(k,\ell),h}+\dim\Lambda_{k-1,h}$, where $\Lambda_{k-1,h}$ is the space of Lagrange multiplier. Numerical results of $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$, $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|$, $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ of the distributional mixed method (39a)-(39b) with $k=1$ and $\ell=0,1$ are shown in Table 1. It is observed that $\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|=\mathcal{O}(h)$, and $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\mathcal{O}(h)$ numerically, which coincide with the theoretical error estimates in (48) and (49). The error $\|{\boldsymbol{u}}-\boldsymbol{u}_{h}\|=\mathcal{O}(h^{\ell+1})$ is also observed but not included in (48) and (49).
Table 1. Errors for the distributional mixed finite element method (39a)-(39b) with $k=1$.
$h$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
# DoFs for $\ell=0$ 2,332 17,240 132,400 1,037,408
# DoFs for $\ell=1$ 2,936 21,424 163,424 1,276,096
$\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$ 1.51E+02 8.62E+01 4.48E+01 2.26E+01
order -- 0.81 0.95 0.99
$\|{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 1.48E+01 5.60E+00 2.39E+00 1.13E+00
order -- 1.40 1.23 1.08
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}(\boldsymbol{u}-\boldsymbol{u}_{h})\|$ 1.46E+02 1.46E+02 1.46E+02 1.46E+02
order -- 0 0 0
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=0$ 1.49E+00 5.23E-01 2.19E-01 1.03E-01
order -- 1.51 1.25 1.09
$\|\boldsymbol{u}-\boldsymbol{u}_{h}\|$ for $\ell=1$ 1.04E+00 2.82E-01 7.32E-02 1.85E-02
order -- 1.88 1.95 1.98
@@ 4.4. Post-processing
Post-processing can be also applied to improve the approximation. It follows from the standard procedure of the stable mixed methods and will be briefly reviewed below.
We will construct a new superconvergent approximation to $\boldsymbol{u}$ in virtue of the optimal result of $\boldsymbol{\sigma}$ in (48). Introduce discrete space
@@
Let $(\boldsymbol{\sigma},0,\boldsymbol{u})$ and $(\boldsymbol{\sigma}_{h},0,\boldsymbol{u}_{h})$ be the solution of the mixed formulation (34a)-(34b) and the mixed finite element method (39a)-(39b) respectively. Assume $\boldsymbol{\sigma}\in H^{k}(\Omega;\mathbb{T})$ and $\boldsymbol{u},{\rm curl\,}\boldsymbol{u}\in H^{k+1}(\Omega;\mathbb{R}^{3})$. Then
@@
$h$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$
# DoFs for $\ell=0$ 2,332 17,240 132,400 1,037,408
# DoFs for $\ell=1$ 2,936 21,424 163,424 1,276,096
$\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 9.25E+00 2.71E+00 7.18E-01 1.83E-01
order -- 1.77 1.92 1.98
$\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|$ 1.52E+02 8.62E+01 4.48E+01 2.26E+01
order -- 0.81 0.95 0.99
Numerical results of the post-processing $\boldsymbol{u}_{h}^{\ast}$ for $k=1$ and $\ell=0,1$ are listed in Table 2. We can observe that $\|{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h^{2})$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}_{\mathcal{T}_{h}}(\boldsymbol{u}-\boldsymbol{u}_{h}^{\ast})\|=\mathcal{O}(h)$ numerically, which are one order higher than $\|{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ and $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|$ respectively. Indeed, $\|{\rm grad\,}_{\mathcal{T}_{h}}{\rm curl\,}({\boldsymbol{u}}-\boldsymbol{u}_{h})\|=\|\boldsymbol{\sigma}-\boldsymbol{\sigma}_{h}\|$.
5. Hybridization of distributional mixed finite element method
In this section we will hybridize the distributional mixed finite element method (39a)-(39b) following the framework in [1]. We introduce a Lagrange multiplier for the tangential-normal continuity of $\boldsymbol{\sigma}$ which can be treat as an approximation of $(\boldsymbol{n}_{F}\times{\rm curl\,}\boldsymbol{u})|_{\mathcal{F}_{h}}$. As $\boldsymbol{\sigma}$ is discontinuous, it can be eliminated element-wise and the size of the resulting linear system is reduced, which is easier to solve than the saddle point system obtained by the mixed method (39a)-(39b). With the hybridized method, we can also establish the equivalence of the mixed method (39a)-(39b) to a weak Galerkin method without stabilization and nonconforming finite element methods in [30, 48].